Write the equation of the line in standard form. Use integer coefficients
y= 3/8x+7/8

Answer by TimothyLamb(4379)

---
y = 3/8x + 7/8
(3/8)x - y = -7/8
---
answer:
3x - 8y = -7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php